Operations Report – March 13th

Crew 190 Operations Report 13th March 2018
SOL: 2
Name of person filing report: Bastien BAIX
Non-nominal systems: –
Notes on non-nominal systems: –

Generator (hours run): Turned off at 10:36 am and turned on at 7:30 pm.
Solar – SOC 84% (Before generator is run at night)
Diesel – ~45%
Propane – 82%
Ethanol Free Gasoline (5 Gallon containers for ATV) – 2 empty jerrycans.

Water (trailer) – 0 gallons.
Water (static) – ~450 gallons
Water (third tank) – 500 gallons.
Trailer to Static Pump used – NO
Water (loft) – Static to Loft Pump used – YES
Water Meter: 132 074.2
Toilet tank emptied: NO

ATV’s Used: 2 & 3
Oil Added: NO
ATV Fuel Used: NO
Hours the ATVs were Used today: 1 hour
Notes on ATVs: –

Deimos rover used: NO
Hours: 116.6
Beginning charge: 100%
Ending charge: –
Currently charging: NO
Notes on Deimos rover: flat tyre

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: NO
Hours: 24.8
Beginning charge: 100%
Ending charge: –
Currently charging: NO

Opportunity rover used: NO
Hours: 20.6
Beginning charge: 100%
Ending charge: –
Currently charging: YES

Curiosity rover used: YES
Hours: 17,2
Beginning charge: 76%
Ending charge: 100%
Currently charging: YES

HabCar used and why, where: –
General notes and comments: –
Summary of internet: ~ 200 Mb remaining
Summary of suits and radios: –
Summary of Hab operations: Ariane & Bastien working on their experiments
Summary of GreenHab operations: Mario & Fred working on their experiments. Aquaponic fish seedlings did not sprout. Suspected due to cold water. Pump turned on and hydroponic solution added to rectify.
Summary of ScienceDome operations: Martin, Sophie, Maximilien &
Michael working on their experiments.
Summary of RAMM operations: –
Summary of health and safety issues: HSO first report submitted

Questions, concerns and requests to Mission Support: –

Greenhab Report – March 13th

Crewmember’s name : Mario Sundic (GreenHab officer & journalist)

Date 13/03/2018

Environmental control:

Ambient with door open.

Shade cloth on.

Fan on, without cooling.

Average temperatures: 30°C

Low : 18°C

High : 37°C

Hours of supplemental light: 0 hours

Changes to crops: No change

Daily water usage for crops: 13 gal

Time(s) of watering for crops: 3

Moringa research observations: ?

Changes to research plants: N/A

Aquaponics: All fish are dead. No seedlings have sprouted. As they were permanently submerged in cold water, preventing them from sprouting, I have put the pump on a 15on/45off timer and added hydroponics solution in the now empty tank. Hopefully this will allow the seeds to sprout and to grow normally in hydroponic conditions, before new fish are brought and the system is perfected.

Support/supplies needed: potting soil

Journalist Report – March 13th

Journalist Report :

This morning, the whole crew managed to wake up at the same time, at a reasonable 8am. This process was helpfully catalysed by Sophie’s uplifting music played on the loudspeaker, and before we knew it, we were downstairs, jumping, squatting and crunching on the first of many morning cereal (which is already starting to run out..), half the crew got ready for an EVA, travelling far to the north workouts to come. After the usual morning, where Max took some soil samples and Bastien made his first (an successful) attempt at drone mapping a patch of land. Meanwhile, the other four members got a head start on their day, beginning work on their experiments. Fred weighed, mixed and potted his Martian soil samples, added HydroGel in half his samples and then planted basil, mint and radishes in them. Michael worked on perfecting his scheduling program, and Max started taking pH measures on soil. On the other hand, Mario made big advancements on his hydroponics project, nearing completion of the first tower, whereas Ariane isolated her first bacteria, worked on her sourdough, and made some more bread prototypes. All in all, our SOL 2 has been the most productive yet, and this has been felt in the crew : the mood is becoming more productive and more exciting as we realise that our projects are finally taking place in the way we wanted them to.

Commander Report – March 13th

Today we started our working day by a gym session given by our crew physicist Sophie. An EVA was planned for 4 of us, Martin, Ariane, Bastien and I, so we just had time to eat breakfast and get dressed with the spacesuits and all material needed for the outing. Everyone was motivated and all procedures were perfectly followed by the EVA’s team. We left at 9.30 a.m. with two ATVs and the Curiosity rover. After a few minutes, we reached “White Moon” on the northern side of the Hab. Our crew engineer Bastien was the EVA team leader and he started mapping the area using his drone. He is currently working on these data in view to create a 3D map. His results seem to be encouraging! During the outing, I collected 3 soil samples for further analysis at the station. Theses samples were collected on “White Moon” and on a hill near “URC North Site”.

We came back at the Hab at 10.30 a.m. after the EVA without unexpected event and we stayed a while in the Hab together having a discussion about missing reports and things that must be improved. We lunched together and then everyone started their experiments. Michael spent his afternoon in the ScienceDome, coding his automatic scheduler. Sophie tried to take pictures of the Sun with the telescope but she is still trying to deal with it.

Ariane began her bacteria culture and the selection of the right strains. Our GreenHab officer almost finished his first hydroponic tower and it will soon be ready for a first test. Regarding other experiments, I spent time with Martin in the ScienceDome analyzing our respective samples. We have already got some results on the soil samples collected in the morning showing a difference of acidity between the samples (“URC North Site” soil is more basic than the others). Frédéric, on his side, added a mix of hydrogel on Martian soil for his mint, basil and radish plantations. Ariane is currently baking bread for our supper and we will prepare dinner. Tonight we will probably play board games to relax and strengthen the team spirit.

Regarding the station, no particular problem was encountered. Tomorrow, we will continue the work on our respective scientific projects and some of us will go on EVA (see EVA request).

EVA Report – March 13th

EVA Report

EVA #3 – 13th March 2018 – Soil 2
Crew members:
Bastien Baix (EVA leader), Ariane Sablon, Maximilien Richald, Martin Roumain

Objectives:
– First drone operations for the cartography project by Bastien.
– Drug samples check by Martin
– Soil samples harvest by Maximilien

Operations:
9:29 – Entering the airlock
9:33 – Leaving the MDRS with 2 ATVs and 1 Rover (Curiosity)
9:45 – Meeting with some local Martian life
9:52 – Arrival at White Moon, samples harvesting, drone cartography and pictures
10:05 – We ride our vehicles further to Beige Moon
10:19 – Leaving Beige Moon
10:27 – Meeting some Martian life again
10:30 – Stop at URC North Site to harvest some red soil samples.
10:35 – Arrival at the Hab
10:36 – Drug samples check, and generator switched off
10: 41 – Entering the airlock
11:05 Leaving the MDRS,

Narrative:

Our third EVA took place on soil 2 at 9:30. I’m Bastien Baix and I was EVA leader today. After our 3 minutes depressurization in the airlock, we were ready to leave the Hab with 3 vehicles. Our destination: White Moon. On our way to this site, we had the surprise to meet some special Martian life that stood on our road (it looked really like a black cow).
Arrived at White Moon, I had the opportunity to finally test my cartography project using the drone. Meanwhile, Maximilien helped by Martin and Ariane were harvesting some soil samples.

After 10 minutes, we decided to go further into Beige Moon and take some pictures and videos with the drone. Indeed, Maximilien had his camera but not his batteries… On our way back, the strange Martian life was still there and we had to get around the specimen. After a last stop at URC North Site for soil samples, we finally arrived at the Hab. Martin checked his drug samples that are located near the solar panels and we also switched off the generator.

This was a long trip but a short EVA in the end.

Vehicles

– Curiosity rover
– ATVs 2 & 3

Astronomy Report – March 13th

Astronomy Report

Name: Sophie Wuyckens Crew: 190
date: 3/13/2018

Sky Conditions: Very clear sky

Wind Conditions: No wind

Observation Start Time: 12:00

Observation End Time: 15:00

Summary: I continued my training with the telescope. I got 2 pictures (see jpeg attached) of the Sun, they are not very fancy but unfortunately the Sun has no real relevant activity. I am trying to create gif and videos of the mini solar eruptions.
Objects Viewed: some little solar eruptions and surface of the Sun

Problems Encountered: No problems

Sol Summary – March 13th

Crew 190 Sol Summary Report
13March2018
Sol 2

Summary Title:
First Rendez-vous with the Martian regolith

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 3, 08:00

Sol Activity Summary:
8:00 Gym
8:20 Breakfast
9:30 EVA: Maximilien, Martin, Ariane, Bastien
9:30 Science/Experimental work: Sophie, Mario, Frederic, Michael
11:00 EVA ended; all crew members switch to SC-work.
13:30 Lunch
14:30 Scientific work: telescope (Sophie), cartography (Bastien), chemistry/biology (Ariane, Martin, Maximilien), GreenHab (Mario, Frederic), operations research (Michael)
18:00 Report redaction
19:00 CapCom

Look Ahead Plan: (automatically generated by our optimizer)
Sol 3:

Sol 3 to 7

Anomalies in work:
None

Weather:

Cloudy in the morning. Clear blue sky in the afternoon, enabling telescope activity. Warm.

Crew Physical Status:

In good shape and still enthusiastic.

EVA:
See EVA report. Mainly soil sampling.

Reports to be file:
Commander report
Journalist report
EVA request for Sol3
Engineer report
HSO report

Crew biographies

(Mission Plan)

Support Requested:
None